OpenStudy (anonymous):

The number of two-digit positive integers for which the units digit is not equal to the tens digit.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi misty =)

OpenStudy (misty1212):

my guess is \(81\) reason as follows: you have nine choices for the tens place digit (it cannot be zero) then another 9 choices for the ones place (it can be zero, but it cannot be the tens digit)

OpenStudy (misty1212):

by the counting principle you get \(9\times 9=81\)

OpenStudy (anonymous):

ouuuuu niceeeee okay got it thanks!

OpenStudy (misty1212):

course i could be wrong maybe @geerky42 has a different answer

OpenStudy (anonymous):

it was right

geerky42 (geerky42):

There are \(99-10+1 = 90\) two digits numbers. Now exclude numbers where tens digit and unit digit are same; 11, 22, 33, 44, ..., 99 There are \(9\) of them. So you have \(90-9 = \boxed{81}\) So @misty1212 is correct.

OpenStudy (anonymous):

hey why do u exclude the 11, 22,33,44 etc.?????

OpenStudy (misty1212):

for which the units digit is not equal to the tens digit.

OpenStudy (anonymous):

but doesnt that just mean 10, 20, 30, 40 etc?

geerky42 (geerky42):

No. "units digit is equal to the tens digit" means 11, 22, 33, etc. So here, we have "units digit is NOT equal to the tens digit", which means 10, 12, 13, ..., 20, 21, 23, ..., 31, 32, 34, ..., etc.

geerky42 (geerky42):

We counted ALL two digits numbers, then we un-counted number where "units digit is equal to the tens digit"